# Enhancing Sum Capacity via Quantum and No-Signaling Cooperation Between Transmitters

### v1

@@ Abstract
Index Terms:
Interference channel, multiple access channel, sum capacity, quantum advantage, non-local game
@@ I Introduction
Quantum information processing has attracted significant attention due to its potential to outperform classical information processing in certain tasks, thus exhibiting quantum advantage. Quantum advantage is not only of theoretical interest but also provides justification for the practical implementation of quantum information processing. Examples of quantum advantage arise in the contexts of algorithmic complexity [1, 2, 3], security [4, 5, 6], privacy [7], non-local games [8, 9, 10], distributed computing [11, 12, 13], and communication over networks [14, 15, 16, 17, 18, 19].
@@
For quantum advantage in communication over networks, the works [14, 15, 16, 18, 17, 19] considered scenarios where multiple transmitters cooperate based on pre-shared resources in transmitting classical messages over classical interference channels (IC) or multiple access channels (MAC). In this scenario, quantum advantage was demonstrated by showing that the sum capacity increases when the transmitters exploit quantum cooperation, as compared to classical cooperation, for specific channels involving pseudo-telepathy games [8, 9, 10]. Beyond quantum cooperation, no-signaling cooperation has also been considered to increase the sum capacity [20, 18, 19, 17].
@@
The aforementioned results were derived by adapting the characteristics of pseudo-telepathy games to the context of network communication. Specifically, a pseudo-telepathy game is a non-local game whose winning probability is strictly less than one with classical cooperation but equal to one with quantum or no-signaling cooperation. Examples of pseudo-telepathy games include the Clauser-Horne-Shimony-Holt (CHSH) game [21], the magic square game [22], and the multiparty parity game [8]. These games exhibit a hierarchy among three types of cooperation: quantum cooperation produces correlations unattainable by classical cooperation, and no-signaling cooperation generates correlations unattainable by quantum cooperation. Interpreted in the context of communication, quantum or no-signaling cooperation between transmitters can generate correlated channel inputs that are not possible through classical cooperation. Consequently, pseudo-telepathy games give rise to network channels in which interference between channel inputs can be mitigated by quantum or no-signaling cooperation, but not by classical cooperation.
@@
The previously considered channels possessing the above properties are as follows. The works [18, 14, 15, 16, 17, 20, 19] defined ICs and MACs where channel inputs are specified as tuples of questions and answers from a pseudo-telepathy game. Additionally, the channels studied in [14, 15, 16, 17, 20] relay the questions from transmitters to one or more receivers noiselessly when the channel inputs satisfy the winning condition of the game. More general channels have also been considered, where the transmitted questions are subject to random erasure [20, 18], or to uniform random noise [19], even when the channel inputs satisfy the winning condition. However, it has not yet been fully determined for which channels the sum capacity can be increased by quantum or no-signaling cooperation.
@@
In this work, we propose a broader class of channels by focusing on the common characteristics shared by the previously studied channels: the interference is weaker when the channel inputs satisfy the winning condition of a pseudo-telepathy game than when they do not. Then, we show the enhancement of the sum capacity through quantum or no-signaling cooperation for such class of channels. Specifically, we propose a class of channels involving pseudo-telepathy games that become less noisy and decompose into parallel weakly symmetric sub-channels when the channel inputs satisfy the winning condition of the game. This class of channels includes the channels studied in prior works [14, 17, 15, 16] as well as new ones not considered in [18, 14, 15, 16, 17, 20, 19].
To show the enhancement, we first derive an upper bound on the sum capacity of the proposed channels, which cannot be achieved by classical cooperation. Next, we show that the sum rate equal to this upper bound can be achieved if the transmitters exploit cooperation that wins the associated pseudo-telepathy game with probability one. Furthermore, we numerically evaluate the amount of enhancement in several examples by applying the generalized Blahut-Arimoto algorithm [23].
@@ II-A Communication over IC without transmitter cooperation
The conventional communication scenario between $K$ transmitter-receiver pairs, $K\geq 2$, over a stationary discrete memoryless interference channel (DM-IC) without feedback has been extensively studied [24]. A DM-IC is characterized as a conditional probability $P_{Y_{[K]}|X_{[K]}}$ from $\mathcal{X}_{[K]}$ to $\mathcal{Y}_{[K]}$, where $\mathcal{X}_{i}$ and $\mathcal{Y}_{i}$ are finite sets denoting an input alphabet at the $i$-th transmitter side and an output alphabet at the $i$-th receiver side, respectively. This channel is stationary and memoryless in the sense that $n$-times use of the channel is given by $P_{Y_{[K]}^{n}|X_{[K]}^{n}}=P_{Y_{[K]}|X_{[K]}}^{\otimes n}$, i.e.,
@@
To date, a single-letter characterization of the capacity region and the sum capacity of a DM-IC is not known in general, even for $K=2$. The best-known inner bound of the capacity region of a general DM-IC is the Han-Kobayashi inner bound [25, 26], and this inner bound is tight for some class of DM-ICs.
@@ II-B Communication over IC with transmitter cooperation
Now, we consider a scenario where transmitters cooperate using some pre-shared resources, which has been studied in [20, 18, 15]. With such cooperation, the transmitters can transmit correlated codewords, and the correlation between them can be useful for mitigating interference, and hence, enlarges the capacity region.
Cooperation between the transmitters is characterized by a conditional probability that can be classified based on the type of pre-shared resource: classical, quantum, or super-quantum resource [10]. Cooperation that utilizes a classical, quantum, or super-quantum resource is termed classical, quantum, or no-signaling cooperation, respectively, and is defined as follows.
@@
Note that a quantum cooperation is not a classical cooperation only if a quantum state $\rho$ in (5) is entangled. Also, there is a strict hierarchy among the above types of cooperations in the containment relation [10], i.e.,
@@
In the case $K=2$, previous works [20, 15, 18] showed that for certain DM-ICs related to pseudo-telepathy games, utilizing quantum or no-signaling cooperation achieves a higher sum capacity than classical cooperation, i.e., $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{NS}}$.
@@ II-C Non-local games
The usefulness of cooperation in a higher hierarchy can be explained in a non-local game scenario [8, 9, 10]. A promise-free $K$-party game $G$ is defined by a tuple of finite sets $(\mathcal{Q}_{[K]},\mathcal{A}_{[K]},\mathcal{W})$ where $\mathcal{W}\subseteq\mathcal{Q}_{[K]}\times\mathcal{A}_{[K]}$. For each $i\in[K]$, $\mathcal{Q}_{i}$ and $\mathcal{A}_{i}$ denote the sets of questions and answers for the $i$-th player, respectively, and $\mathcal{W}$ denotes the winning condition. [^1] A promise is a subset $P$ of the set of questions $\mathcal{Q}_{[K]}$. For a $K$-party game $G$ with promise $P$, a referee selects questions $Q_{[K]}$ only from $P$. In this paper, we only consider a promise-free game, i.e., $P=\mathcal{Q}_{[K]}$.
@@
There are games that show strict gaps between the maximum winning probabilities that can be achieved by different types of cooperations. One of the most representative examples is the Clauser–Horne–Shimony–Holt (CHSH) game [21]. The CHSH game $G_{\mathrm{CHSH}}$ is a 2-party game such that
@@
One example of a $\mathrm{Q}$-pseudo-telepathy game is the magic square game $G_{\mathrm{MS}}$ [22]. It is a two-party game that consists of
@@
Beyond two-party $\mathrm{Q}$-pseudo-telepathy game, there are $K$-party $\mathrm{Q}$-pseudo-telepathy games for all $K\geq 3$, known as the (promise-free version of) $K$-party parity ($K$-$P$) game [8, 17]. The $K$-$P$ game $G_{K\mathrm{-PP}}$ consists of
@@ III Main results
Our main result is on showing the enhancement of sum capacity according to the hierarchy of cooperation, i.e., $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{NS}}$, for a certain class of DM-ICs and DM-MACs. The proposed class includes all the channels previously used to demonstrate such separation in previous work [14, 17, 16, 15], as well as many other channels beyond the channels considered in [18, 14, 15, 16, 17, 20, 19].
@@ III-A DM-ICs from non-local games
Inspired by $K$-party games and the previous works [18, 14, 15, 16, 17, 20, 19], we propose a class of DM-ICs for which quantum or no-signaling cooperation can achieve strictly greater sum capacity than classical cooperation. The basic idea is to design a DM-IC $P_{Y_{[K]}|X_{[K]}}$ that is less noisy when channel inputs $x_{[K]}$ satisfy the winning condition of a game $G$ than when they do not satisfy. Specifically, the conditional entropy $H(Y_{[K]}|X_{[K]}=x_{[K]})$ is smaller when $x_{[K]}\in\mathcal{W}$ than when $x_{[K]}\notin\mathcal{W}$. Further, we set channel input $x_{i}$ to a pair of question and answer $(q_{i},a_{i})$ for $G$, and design $P_{Y_{[K]}|X_{[K]}}$ such that it decomposes into parallel sub-channels, i.e., $P_{Y_{[K]}|X_{[K]}}=\prod_{i\in[K]}P_{Y_{i}|Q_{i}}$, whenever $x_{[K]}\in\mathcal{W}$. Here, $a_{[K]}$ is discarded if the winning condition is satisfied. By requiring each sub-channel $P_{Y_{i}|Q_{i}}$ to be weakly symmetric [27, Sec. 7.2], we can show the enhancements of sum capacity for some of these channels. The details are as follows.
@@
with $0\leq\eta_{l}<\eta_{w}\leq 1$. These are instances of $G$-game channels that are not considered in the previous works [18, 14, 15, 16, 17, 20, 19] (together with the identification of DM-ICs as DM-ICs mentioned in Remark 1). The channels considered in the previous works [14, 17, 16, 15] to show the enhancements of sum capacity are instances of a game channel of this form with $\eta_{w}=1$ and $\eta_{l}=0$.
@@
Previous work [19] considered a class of DM-MACs $P_{\tilde{Y}|X_{[K]}}$ determined by a game $G$, such that $\mathcal{X}_{i}=(\mathcal{Q}_{i},\mathcal{A}_{i})$, $\tilde{\mathcal{Y}}=\mathcal{Q}_{[K]}$, and

### v2

@@ I Introduction
Quantum information processing has attracted significant attention due to its potential to outperform classical information processing in certain tasks, thus exhibiting quantum advantage. Quantum advantage is not only of theoretical interest but also provides justification for the practical implementation of quantum information processing. Examples of quantum advantage arise in the contexts of algorithmic complexity [shorPolynomialTimeAlgorithmsPrime1997, groverFastQuantumMechanical1996, deutsch_rapid_1992], security [bennettQuantumCryptographyPublic2014, ekertQuantumCrypt1991, bennett_quantum_1992], privacy [namQuantumAdvantagePrivate2025], non-local games [brassardQuantumPseudoTelepathy2005a, cleveConsequencesLimitsNonlocal2004, brunnerBellNonlocality2014], distributed computing [cleve_substituting_1997, bar-yossef_exponential_2004, buhrmanNonlocalityCommunicationComplexity2010a], and communication over networks [leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, YunRaiBae2020, yunNonLocalQuantumAdvantages2023].
@@
For quantum advantage in communication over networks, the works [leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, YunRaiBae2020, seshadriSeparationCorrelationAssistedSum2023, yunNonLocalQuantumAdvantages2023] considered scenarios where multiple transmitters cooperate based on pre-shared resources in transmitting classical messages over classical interference channels (IC) or multiple access channels (MAC). In this scenario, quantum advantage was demonstrated by showing that the sum capacity increases when the transmitters exploit quantum cooperation, as compared to classical cooperation, for specific channels involving pseudo-telepathy games [brassardQuantumPseudoTelepathy2005a, cleveConsequencesLimitsNonlocal2004, brunnerBellNonlocality2014]. Beyond quantum cooperation, no-signaling cooperation has also been considered to increase the sum capacity [quekQuantumSuperquantumEnhancements2017, YunRaiBae2020, yunNonLocalQuantumAdvantages2023, seshadriSeparationCorrelationAssistedSum2023]. While there have also been studies exploring more general scenarios involving no-signaling cooperation among both transmitters and receivers, as well as network channel models beyond MACs and ICs [fawziBroadcastChannelCoding2024, fawziMultipleAccessChannelCoding2024, notzelEntanglementEnabledCommunication2020], we focus on MACs and ICs with quantum and no-signaling cooperation between transmitters, as considered in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023]. Even in these simpler settings, it remains an open question under which classes of channels such an improvement can be achieved.
@@
In the aforementioned settings, previous results in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023] were derived by adapting the characteristics of pseudo-telepathy games to the context of network communication. Specifically, a pseudo-telepathy game is a non-local game whose winning probability is strictly less than one with classical cooperation but equal to one with quantum or no-signaling cooperation. Examples of pseudo-telepathy games include the Clauser-Horne-Shimony-Holt (CHSH) game [clauserProposedExperimentTest1969], the magic square game [aravindBellsTheoremInequalities2002], and the multiparty parity game [brassardQuantumPseudoTelepathy2005a]. These games exhibit a hierarchy among three types of cooperation: quantum cooperation produces correlations unattainable by classical cooperation, and no-signaling cooperation generates correlations unattainable by quantum cooperation. Interpreted in the context of communication, quantum or no-signaling cooperation between transmitters can generate correlated channel inputs that are not possible through classical cooperation. Consequently, pseudo-telepathy games give rise to network channels in which interference between channel inputs can be mitigated by quantum or no-signaling cooperation, but not by classical cooperation.
@@
The previously considered channels possessing the above properties are as follows. The works [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023] defined ICs and MACs where channel inputs are specified as tuples of questions and answers from a pseudo-telepathy game. Additionally, the channels studied in [leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017] relay the questions from transmitters to one or more receivers noiselessly when the channel inputs satisfy the winning condition of the game. More general channels have also been considered, where the transmitted questions are subject to random erasure [quekQuantumSuperquantumEnhancements2017, YunRaiBae2020], or to uniform random noise [yunNonLocalQuantumAdvantages2023], even when the channel inputs satisfy the winning condition. However, it has not yet been fully determined for which channels the sum capacity can be increased by quantum or no-signaling cooperation.
@@
In this work, we propose a broader class of channels by focusing on the common characteristics shared by the previously studied channels: the interference is weaker when the channel inputs satisfy the winning condition of a pseudo-telepathy game than when they do not. Then, we show the enhancement of the sum capacity through quantum or no-signaling cooperation for such class of channels. Specifically, we propose a class of channels involving pseudo-telepathy games that become less noisy and decompose into parallel weakly symmetric sub-channels when the channel inputs satisfy the winning condition of the game. This class of channels includes the channels studied in prior works [leditzkyPlayingGamesMultiple2020, seshadriSeparationCorrelationAssistedSum2023, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025] as well as new ones not considered in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023]. To show the enhancement, we first derive an upper bound on the sum capacity of the proposed channels, which cannot be achieved by classical cooperation. Next, we show that the sum rate equal to this upper bound can be achieved if the transmitters exploit cooperation that wins the associated pseudo-telepathy game with probability one.
@@ II-A Communication over IC without transmitter cooperation
The conventional communication scenario between $K$ transmitter-receiver pairs, $K\geq 2$, over a stationary discrete memoryless interference channel (DM-IC) without feedback has been extensively studied [elgamalNetworkInformationTheory2011]. A DM-IC is characterized as a conditional probability $P_{Y_{[K]}|X_{[K]}}$ from $\mathcal{X}_{[K]}$ to $\mathcal{Y}_{[K]}$, where $\mathcal{X}_{i}$ and $\mathcal{Y}_{i}$ are finite sets denoting an input alphabet at the $i$-th transmitter side and an output alphabet at the $i$-th receiver side, respectively. This channel is stationary and memoryless in the sense that $n$-times use of the channel is given by $P_{Y_{[K]}^{n}|X_{[K]}^{n}}=P_{Y_{[K]}|X_{[K]}}^{\otimes n}$, i.e.,
@@
To date, a single-letter characterization of the capacity region and the sum capacity of a DM-IC is not known in general, even for $K=2$. The best-known inner bound of the capacity region of a general DM-IC is the Han-Kobayashi inner bound [hanNewAchievableRate1981, chongHanKobayashiRegion2008], and this inner bound is tight for some class of DM-ICs.
@@ II-B Communication over IC with transmitter cooperation
Now, we consider a scenario where transmitters cooperate using some pre-shared resources, which has been studied in [quekQuantumSuperquantumEnhancements2017, YunRaiBae2020, hawellekInterferenceChannelEntangled2025]. With such cooperation, the transmitters can transmit correlated codewords, and the correlation between them can be useful for mitigating interference, and hence, enlarges the capacity region.
Cooperation between the transmitters is characterized by a conditional probability that can be classified based on the type of pre-shared resource: classical, quantum, or super-quantum resource [brunnerBellNonlocality2014]. Cooperation that utilizes a classical, quantum, or super-quantum resource is termed classical, quantum, or no-signaling cooperation, respectively, and is defined as follows.
@@
Note that a quantum cooperation is not a classical cooperation only if a quantum state $\rho$ in (5) is entangled. Also, there is a strict hierarchy among the above types of cooperations in the containment relation [brunnerBellNonlocality2014], i.e.,
@@
In the case $K=2$, previous works [quekQuantumSuperquantumEnhancements2017, hawellekInterferenceChannelEntangled2025, YunRaiBae2020] showed that for certain DM-ICs related to pseudo-telepathy games, utilizing quantum or no-signaling cooperation achieves a higher sum capacity than classical cooperation, i.e., $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{NS}}$.
@@ II-C Non-local games
The usefulness of cooperation in a higher hierarchy can be explained in a non-local game scenario [brassardQuantumPseudoTelepathy2005a, cleveConsequencesLimitsNonlocal2004, brunnerBellNonlocality2014]. A promise-free $K$-party game $G$ is defined by a tuple of finite sets $(\mathcal{Q}_{[K]},\mathcal{A}_{[K]},\mathcal{W})$ where $\mathcal{W}\subseteq\mathcal{Q}_{[K]}\times\mathcal{A}_{[K]}$. For each $i\in[K]$, $\mathcal{Q}_{i}$ and $\mathcal{A}_{i}$ denote the sets of questions and answers for the $i$-th player, respectively, and $\mathcal{W}$ denotes the winning condition. [^1] A promise is a subset $P$ of the set of questions $\mathcal{Q}_{[K]}$. For a $K$-party game $G$ with promise $P$, a referee selects questions $Q_{[K]}$ only from $P$. In this paper, we only consider a promise-free game, i.e., $P=\mathcal{Q}_{[K]}$.
@@
There are games that show strict gaps between the maximum winning probabilities that can be achieved by different types of cooperations. One of the most representative examples is the Clauser–Horne–Shimony–Holt (CHSH) game [clauserProposedExperimentTest1969]. The CHSH game $G_{\mathrm{CHSH}}$ is a 2-party game such that
@@
One example of a $\mathrm{Q}$-pseudo-telepathy game is the magic square game $G_{\mathrm{MS}}$ [aravindBellsTheoremInequalities2002]. It is a two-party game that consists of
@@
Beyond two-party $\mathrm{Q}$-pseudo-telepathy game, there are $K$-party $\mathrm{Q}$-pseudo-telepathy games for all $K\geq 3$, known as the (promise-free version of) $K$-party parity ($K$-$P$) game [brassardQuantumPseudoTelepathy2005a, seshadriSeparationCorrelationAssistedSum2023]. The $K$-$P$ game $G_{K\mathrm{-PP}}$ consists of
@@ III Main results
Our main result is on showing the enhancement of sum capacity according to the hierarchy of cooperation, i.e., $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{Q}}$ or $\mathcal{C}_{s}^{\mathrm{C}}<\mathcal{C}_{s}^{\mathrm{NS}}$, for a certain class of DM-ICs and DM-MACs. The proposed class includes all the channels previously used to demonstrate such separation in previous work [leditzkyPlayingGamesMultiple2020, seshadriSeparationCorrelationAssistedSum2023, peregMultipleAccessChannelEntangled2025, hawellekInterferenceChannelEntangled2025], as well as many other channels beyond the channels considered in [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023].
@@ III-A DM-ICs from non-local games
Inspired by $K$-party games and the previous works [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023], we propose a class of DM-ICs for which quantum or no-signaling cooperation can achieve strictly greater sum capacity than classical cooperation. The basic idea is to design a DM-IC $P_{Y_{[K]}|X_{[K]}}$ that is less noisy when channel inputs $x_{[K]}$ satisfy the winning condition of a game $G$ than when they do not satisfy. Specifically, the conditional entropy $H(Y_{[K]}|X_{[K]}=x_{[K]})$ is smaller when $x_{[K]}\in\mathcal{W}$ than when $x_{[K]}\notin\mathcal{W}$. Further, we set channel input $x_{i}$ to a pair of question and answer $(q_{i},a_{i})$ for $G$, and design $P_{Y_{[K]}|X_{[K]}}$ such that it decomposes into parallel sub-channels, i.e., $P_{Y_{[K]}|X_{[K]}}=\prod_{i\in[K]}P_{Y_{i}|Q_{i}}$, whenever $x_{[K]}\in\mathcal{W}$. Here, $a_{[K]}$ is discarded if the winning condition is satisfied. By requiring each sub-channel $P_{Y_{i}|Q_{i}}$ to be weakly symmetric [coverElementsInformationTheory2012, Sec. 7.2], we can show the enhancements of sum capacity for some of these channels. The details are as follows.
@@
with $0\leq\eta_{l}<\eta_{w}\leq 1$. These are instances of $G$-game channels that are not considered in the previous works [YunRaiBae2020, leditzkyPlayingGamesMultiple2020, hawellekInterferenceChannelEntangled2025, peregMultipleAccessChannelEntangled2025, seshadriSeparationCorrelationAssistedSum2023, quekQuantumSuperquantumEnhancements2017, yunNonLocalQuantumAdvantages2023] (together with the identification of DM-ICs as DM-ICs mentioned in Remark 1). The channels considered in the previous works [leditzkyPlayingGamesMultiple2020, seshadriSeparationCorrelationAssistedSum2023, peregMultipleAccessChannelEntangled2025, hawellekInterferenceChannelEntangled2025] to show the enhancements of sum capacity are instances of a game channel of this form with $\eta_{w}=1$ and $\eta_{l}=0$.
@@
Previous work [yunNonLocalQuantumAdvantages2023] considered a class of DM-MACs $P_{\tilde{Y}|X_{[K]}}$ determined by a game $G$, such that $\mathcal{X}_{i}=(\mathcal{Q}_{i},\mathcal{A}_{i})$, $\tilde{\mathcal{Y}}=\mathcal{Q}_{[K]}$, and